# Multi-Modal Federated Learning for Cancer Staging over Non-IID Datasets with Unbalanced Modalities

### v1

@@ V-D Hyperparamters Tuning
The major hyperparameters tuned in our simulations are the number of layers for the mRNA modality’s encoder depth, the initial learning-rate and its decay-rate. For the mRNA modality’s encoder’s depth, we conducted three experiments with three neural network architectures with multiple layer depths which are listed in Table II. For the initial learning-rate, we chose five values to start the ML training with $\{10^{-2},10^{-3},10^{-4},10^{-5},10^{-6}\}$. The decay-rates for the learning-rate were searched between the values $\{0.99,0.95,0.90,0.85\}$. The best performance for different mixtures of these parameters was observed for the first $50$ global FL steps and the medium depth, $10^{-4}$ initial learning rate, and $0.99$ decay rate over each global round were chosen. We used $K=20$ local SGD iterations on each institution’s ML model. Also, dividing the datasets mentioned in Sec. V-A, we simulated $9$ institutions with different subsets of modalities.
TABLE II: Different architectures for the mRNA encoder.
Layer Type
Model Depth
Small
Medium
Large
Input Layer
FC (20531, ReLU)
Layer 1
FC (4096, ReLU)
FC (8192, ReLU)
FC (16384, ReLU)
Layer 2
FC (2048, ReLU)
FC (4096, ReLU)
FC (8192, ReLU)
Layer 3
FC (512, ReLU)
FC (2048, ReLU)
FC (4096, ReLU)
Layer 4
FC (128, ReLU)
FC (512, ReLU)
FC (2048, ReLU)
Layer 5
FC (64, ReLU)
FC (128, ReLU)
FC (512, ReLU)
Layer 6
FC (32, ReLU)
FC (64, ReLU)
FC (128, ReLU)
Layer 7
FC (32, ReLU)
FC (64, ReLU)
Layer 8
FC (32, ReLU)
Output Layer
FC (2, LogSoftmax)
@@ V-E1 ML training under different learning methods
To test the performance of our method against the baselines, we devised an experiment trying to simulate the methods under the same circumstances. In particular, all of the ML models start from the same initial point/model in order to avoid having random initial weights affect the performance of the models. We took the datasets mentioned in Sec. V-A and created different subsets from each of them to account for the multiple institutions that we have in our network. We acquired the results gathered in Table III for the evolution of performance of each method over time. As can be seen, CM-FL + DGB already surpasses the CM-FL method. However, as the training proceeds, using our entire methodology with the addition of PCW, we can see a better performance. This reveals that the client weighting method used in PCW, although it can slow down the training at the beginning, can effectively circumvent the local ML model bias, especially as ML training proceeds. The table further shows that, due to the unique challenges introduced by the unbalanced distribution of data modalities and labels, baselines that are not designed to handle such non-uniformaties (CM-FL, HGB-FL) fail to train the model effectively.
TABLE III: Accuracy of different methods throughout the global aggregations.
Global Round Method
CM-FL
HGB-FL
CM-FL + DGB
Our Method
UM-FL (Upperbound)
1
$\simeq$ 61.80
$\simeq$ 48.72
$\simeq$ 63.49
$\simeq$ 39.32
$\simeq$ 51.34
100
$\simeq$ 62.92
$\simeq$ 67.52
$\simeq$ 63.34
$\simeq$ 61.80
$\simeq$ 50.89
200
$\simeq$ 59.55
$\simeq$ 65.81
$\simeq$ 63.64
$\simeq$ 64.04
$\simeq$ 51.34
300
$\simeq$ 62.92
$\simeq$ 63.25
$\simeq$ 64.24
$\simeq$ 68.54
$\simeq$ 66.52
400
$\simeq$ 61.80
$\simeq$ 62.39
$\simeq$ 66.43
$\simeq$ 71.91
$\simeq$ 79.46
500
$\simeq$ 64.04
$\simeq$ 64.10
$\simeq$ 67.13
$\simeq$ 74.16
$\simeq$ 79.02
600
$\simeq$ 61.80
$\simeq$ 62.39
$\simeq$ 69.21
$\simeq$ 74.16
$\simeq$ 79.02
@@ V-E2 ML training under different levels of non-uniformity across the modalities
Since we are among the first to study an FL system with an unbalanced number of modalities across clients/institutions, we study the impact of varying the non-uniformity of the number of modalities across institutions. We compare the performance of our method against the baselines under different levels of non-uniformity to show the robustness of our method against the distribution of modalities across institutions. Since we have three modalities, we can have either uni-modal clients, bi-modal clients, or tri-modal clients. We conduct our comparisons in three settings where most institutions possess the dominant number of modalities:
@@
We compare the performance (in terms of prediction accuracy) of our method against the baselines in Table IV. It can be seen that our method outperforms the other baselines by a notable margin. Also, reducing the number of modalities per client (i.e., having all tri-modal clients vs. having all uni-modal clients) leaves a marginal impact on the performance of our method (i.e., $1.33\%$) as compared to baselines, being $1.76\%$ and $2.13\%$. This implies the robustness of our method against the level of heterogeneity of the modalities across institutions.
TABLE IV: Global model accuracy for each dominant modality scenario.
Method Dominant Modality
Uni-Modal
Bi-Modal
Tri-Modal
CM-FL
$\simeq$ 60.12
$\simeq$ 60.56
$\simeq$ 62.25
CM-FL + DGB
$\simeq$ 69.44
$\simeq$ 70.13
$\simeq$ 71.20
Our Method
$\simeq$ 74.30
$\simeq$ 74.83
$\simeq$ 75.63
V-E3 ML training under different number of modalities involved
In addition to heterogeneity of number of modalities across the system, the quality of data of each modality is different across institutions (i.e., some modalities may contribute more to the overall performance of the ML model as compared to the rest). To investigate this, we train the ML models and gather the accuracy of the global model under availability of different subsets of modalities across institutions in Table V. The results show that our method outperforms the baselines under various data qualities. Also, the clinical data shows the lowest quality data as it leads to the lowest performance when combined with other modalities, while the mRNA modality has the highest quality data. This observation can further open the door to research on multi-modal FL in the presence of different quality data and attention mechanisms, which can put non-uniform emphasis on different types of data.
@@ V-E3 ML training under different number of modalities involved
TABLE V: Accuracy of different methods under different subsets of modalities.
Method Modality Subset
mRNA+Image + Clinical
mRNA + Image
mRNA + Clinical
Image + Clinical
CM-FL
$\simeq$ 61.80
$\simeq$ 61.25
$\simeq$ 55.32
$\simeq$ 50.43
CM-FL + DGB
$\simeq$ 69.21
$\simeq$ 65.23
$\simeq$ 55.81
$\simeq$ 54.43
Our Method
$\simeq$ 74.16
$\simeq$ 68.42
$\simeq$ 58.12
$\simeq$ 55.84

### v2

@@ V-D Hyperparamters Tuning
The major hyperparameters tuned in our simulations are the number of layers for the mRNA modality’s encoder depth, the initial learning-rate and its decay-rate. For the mRNA modality’s encoder’s depth, we conducted three experiments with three neural network architectures with multiple layer depths which are listed in Table VI. For the initial learning-rate, we chose five values to start the ML training with $\{10^{-2},10^{-3},10^{-4},10^{-5},10^{-6}\}$. The decay-rates for the learning-rate were searched between the values $\{0.99,0.95,0.90,0.85\}$. The best performance for different mixtures of these parameters was observed for the first $50$ global FL steps and the medium depth, $10^{-4}$ initial learning rate, and $0.99$ decay rate over each global round were chosen. We used $K=20$ local SGD iterations on each institution’s ML model. Also, dividing the datasets mentioned in Sec. V-A, we simulated $9$ institutions with different subsets of modalities.
(a) Binary Classification, Type-Based Heterogeneity
(b) Multi-Class Classification, Type-Based Heterogeneity
(c) Binary Classification, Class-Based Heterogeneity
(d) Multi-Class Classification, Type-Based Heterogeneity
Fig. 5: Bar plot of the values of $\overline{\rho}_{C,n}^{(t)}$ (y-axis) for all institutions $n\in\mathcal{N}$ (x-axis) for each scenario: (a) binary classification, type-based, (b) binary classification, class-based, (c) multi-class classification, type-based, and (d) multi-class classification, class-based. The numbers beside each bar plot indicate the category of the institution’s data based on Table IV for (a) and (c) and Table V for (b) and (d).
@@ V-E1 ML training under different learning methods
To test the performance of our method against the baselines, we devised an experiment trying to simulate the methods under the same circumstances. In particular, all of the ML models start from the same initial point/model in order to avoid having random initial weights affecting the performance of the models. For each heterogeneity case (i.e., type-based and class-based), we used the simulated institutions with the distributions in Table IV and V. The results in Tables VII and VIII show the evolution of performance of each method over time in both binary and multi-class cases for type-based and class-based heterogeneity scenarios, respectively. Focusing on Table VII, it is shown that the addition of DGB initially results in a better performance, from the other methods. However, since it does not explicitly take into account the heterogeneity of data across modality combinations, it fails to retain improvement in global model performance in the later stages of training. This reveals that the client weighting method used in PCW can effectively circumvent the local ML model bias caused by non-iid data across the institutions, especially as the ML training proceeds. This observation is consistent for both binary and multi-class classification scenarios presented in the table, where our method reaches the closest performance to that of the upper-bound (i.e., UM-FL) in the later stages of training. The table further shows that, due to the unique challenges introduced by the unbalanced distribution of data modalities and labels, baselines that are not designed to handle such non-uniformities (CM-FL, HGB-FL) fail to train the model effectively: CM-FL saturates early and starts showing overfitting behavior, while HGB-FL exhibits the same behavior later in the training. The same behavior is observed in Table VIII, where class-based heterogeneity is present across institutions’ local datasets. Comparing the results between Table VII and VIII, it can be seen that the methods show a slightly lower performance under type-based heterogeneity (i.e., Table VII) compared to class-based heterogeneity (i.e., Table VIII). This is due to the fact that, as discussed in Sec. V-A, type-based heterogeneity is a more fundamental level of data heterogeneity, imposing which will lead to more pronounced impact/drop on the performance.
@@ V-E2 ML training under different levels of non-uniformity across the modalities
Since we are among the first to study an FL system with an unbalanced number of modalities across clients/institutions, we study the impact of varying the non-uniformity of the number of modalities across institutions. We compare the performance of our method against the baselines under different levels of non-uniformity to show the robustness of our method against the distribution of modalities across institutions. Since we have three modalities, we can have either uni-modal institutions, bi-modal institutions, or tri-modal institutions. We conduct our comparisons in three settings where most institutions possess the dominant number of modalities:
@@
TABLE XI: Accuracy (%) of different methods under different subsets of modalities for both binary classification and multi-class classification cases in the presence of type-based heterogeneity.
Binary Classification Multi-Class Classification
Method Subset {mRNA, Image, Clinical} {mRNA, Image} {mRNA, Clinical} {Image, Clinical} Method Subset {mRNA, Image, Clinical} {mRNA, Image} {mRNA, Clinical} {Image, Clinical}
CM-FL $\simeq$ 61.80 $\simeq$ 61.25 $\simeq$ 55.32 $\simeq$ 50.43 CM-FL $\simeq$ 48.12 $\simeq$ 47.72 $\simeq$ 46.48 $\simeq$ 46.25
CM-FL + DGB $\simeq$ 69.21 $\simeq$ 65.23 $\simeq$ 55.81 $\simeq$ 54.43 CM-FL + DGB $\simeq$ 53.83 $\simeq$ 50.18 $\simeq$ 48.43 $\simeq$ 47.69
Our Method $\simeq$ 74.16 $\simeq$ 68.42 $\simeq$ 58.12 $\simeq$ 55.84 Our Method $\simeq$ 67.84 $\simeq$ 66.65 $\simeq$ 60.72 $\simeq$ 60.44
Centralized (Upperbound) $\simeq$ 87.64 $\simeq$ 80.43 $\simeq$ 74.13 $\simeq$ 72.81 Centralized (Upperbound) $\simeq$ 80.39 $\simeq$ 76.84 $\simeq$ 75.26 $\simeq$ 74.32
TABLE XII: Accuracy (%) of different methods under different subsets of modalities for both binary classification and multi-class classification cases in the presence of class-based heterogeneity.
Binary Classification Multi-Class Classification
Method Subset {mRNA, Image, Clinical} {mRNA, Image} {mRNA, Clinical} {Image, Clinical} Method Subset {mRNA, Image, Clinical} {mRNA, Image} {mRNA, Clinical} {Image, Clinical}
CM-FL $\simeq$ 65.24 $\simeq$ 63.77 $\simeq$ 53.22 $\simeq$ 52.14 CM-FL $\simeq$ 49.49 $\simeq$ 47.98 $\simeq$ 46.97 $\simeq$ 46.46
CM-FL + DGB $\simeq$ 71.13 $\simeq$ 69.82 $\simeq$ 56.12 $\simeq$ 54.82 CM-FL + DGB $\simeq$ 57.07 $\simeq$ 54.04 $\simeq$ 49.49 $\simeq$ 46.97
Our Method $\simeq$ 76.31 $\simeq$ 70.21 $\simeq$ 60.14 $\simeq$ 57.17 Our Method $\simeq$ 69.92 $\simeq$ 66.75 $\simeq$ 61.42 $\simeq$ 60.74
Centralized (Upperbound) $\simeq$ 87.64 $\simeq$ 80.43 $\simeq$ 74.13 $\simeq$ 72.81 Centralized (Upperbound) $\simeq$ 80.39 $\simeq$ 76.84 $\simeq$ 75.26 $\simeq$ 74.32
We conduct experiments for both binary and multi-class classification cases, considering both type-based and class-based heterogeneities. The results are presented in Tables IX, and X. Focusing on type-based heterogeneity (i.e., Table IX), it can be seen that our method outperforms the other baselines by a notable margin. Also, reducing the number of modalities per institution (i.e., majority of tri-modal institutions vs. majority of uni-modal institutions) leaves a marginal impact on the performance of our method (i.e., $2.07\%$ in binary case and $2.74\%$ in multi-class case) as compared to baselines (being $2.13\%$ and $3.69\%$ in binary case and $2.26\%$ and $3.98\%$ in multi-class case). This implies the robustness of our method against the level of heterogeneity of the modalities across institutions. The same deduction can be made from inspecting the results in Table X, focusing on the class-based heterogeneity. Similar to the observation made in Sec. V-E1, comparing the results in Table IX and X reveals a slight decrease in the performance under type-based heterogeneity which can be explained by the more pronounced impact of the type-based heterogeneity on ML model training as discussed in Sec. V-A.
V-E3 ML training under different contributing of modalities involved
In addition to heterogeneity of number of modalities across the system, the richness of data of each modality is different across institutions (i.e., some modalities may contribute more to the overall performance of the ML model as compared to the rest). To investigate this, we train the ML models and gather the accuracy of the global model under availability of different subsets of modalities across institutions. We show the results of these experiments conducted under type-based heterogeneity in Table XI and under class-based heterogeneity in Table XII. The results in both tables under both binary and multi-class classification cases show that our method outperforms the baselines under various data modality richness levels. Also, the clinical data shows the lowest quality data as it leads to the lowest performance when combined with other modalities, while the mRNA modality has the highest quality data. This observation can further open the door to research on multi-modal FL in the presence of different quality data and attention mechanisms, which can put non-uniform emphasis on different types of data.
TABLE XIII: Final global model accuracy (%) across various cohorts’ data for models trained under type-based heterogeneity.
Binary Classification Multi-Class Classification
Method Cohort BRCA LUSC LIHC Method Cohort BRCA LUSC LIHC
CM-FL $\simeq$ 75.71 $\simeq$ 60.71 $\simeq$ 56.00 CM-FL $\simeq$ 54.78 $\simeq$ 46.34 $\simeq$ 43.90
HGB-FL $\simeq$ 71.42 $\simeq$ 67.86 $\simeq$ 68.00 HGB-FL $\simeq$ 72.85 $\simeq$ 53.66 $\simeq$ 56.10
CM-FL + DGB $\simeq$ 71.43 $\simeq$ 68.57 $\simeq$ 68.00 CM-FL + DGB $\simeq$ 61.21 $\simeq$ 51.22 $\simeq$ 48.78
Our Method $\simeq$ 75.71 $\simeq$ 75.00 $\simeq$ 72.00 Our Method $\simeq$ 68.96 $\simeq$ 68.29 $\simeq$ 65.85
TABLE XIV: Final global model accuracy (%) across various cohorts’ data for models trained under class-based heterogeneity.
Binary Classification Multi-Class Classification
Method Cohort BRCA LUSC LIHC Method Cohort BRCA LUSC LIHC
CM-FL $\simeq$ 71.43 $\simeq$ 64.29 $\simeq$ 60.00 CM-FL $\simeq$ 53.45 $\simeq$ 48.78 $\simeq$ 46.34
HGB-FL $\simeq$ 72.36 $\simeq$ 67.85 $\simeq$ 68.00 HGB-FL $\simeq$ 70.69 $\simeq$ 58.53 $\simeq$ 60.98
CM-FL + DGB $\simeq$ 77.14 $\simeq$ 67.86 $\simeq$ 68.00 CM-FL + DGB $\simeq$ 61.21 $\simeq$ 53.66 $\simeq$ 56.10
Our Method $\simeq$ 80.00 $\simeq$ 78.57 $\simeq$ 72.00 Our Method $\simeq$ 70.69 $\simeq$ 70.73 $\simeq$ 68.29
V-E4 Comparing the values of similarity metric $\overline{\rho}_{C,n}^{(t)}$ during training for different levels of heterogeneity
In order to demonstrate how our proposed PCW method captures and reflects the heterogeneity of data across institutions during the learning process, we provide the bar plot of the values of the similarity metric $\overline{\rho}_{C,n}^{(t)}$ given by (19) – used in obtaining the adjusted average train and validation loss across the institutions in (20) – during the model training period.
TABLE XV: Comparison of iid vs non-iid setting accuracy (%) throughout the global aggregations under binary classification for type-based and class-based heterogeneity cases.
iid Data     non-iid Data, Type-based     non-iid Data, Class-based
Global Round Method
CM - FL + DGB CM - FL + DGB + PCW     CM - FL + DGB CM - FL + DGB + PCW     CM - FL + DGB CM - FL + DGB + PCW Centralized + GB (Upperbound)
0 $\simeq$ 52.34 $\simeq$ 52.34     $\simeq$ 52.34 $\simeq$ 52.34     $\simeq$ 50.16 $\simeq$ 50.16 $\simeq$ 55.14
100 $\simeq$ 65.69 $\simeq$ 65.24     $\simeq$ 63.34 $\simeq$ 61.80     $\simeq$ 65.46 $\simeq$ 63.48 $\simeq$ 68.37
200 $\simeq$ 68.54 $\simeq$ 67.82     $\simeq$ 63.64 $\simeq$ 64.04     $\simeq$ 67.18 $\simeq$ 66.80 $\simeq$ 72.49
300 $\simeq$ 71.77 $\simeq$ 70.16     $\simeq$ 64.24 $\simeq$ 68.54     $\simeq$ 68.12 $\simeq$ 68.23 $\simeq$ 76.30
400 $\simeq$ 72.58 $\simeq$ 74.19     $\simeq$ 66.43 $\simeq$ 71.91     $\simeq$ 69.57 $\simeq$ 73.28 $\simeq$ 81.45
500 $\simeq$ 75.80 $\simeq$ 76.61     $\simeq$ 67.13 $\simeq$ 74.16     $\simeq$ 70.43 $\simeq$ 76.24 $\simeq$ 83.87
600 $\simeq$ 78.23 $\simeq$ 79.03     $\simeq$ 69.21 $\simeq$ 74.16     $\simeq$ 71.13 $\simeq$ 76.31 $\simeq$ 85.48
TABLE XVI: Comparison of iid vs non-iid setting accuracy (%) throughout the global aggregations under multi-class classification for type-based and class-based heterogeneity cases.
iid Data     non-iid Data, Type-based     non-iid Data, Class-based
Global Round Method
CM - FL + DGB CM - FL + DGB + PCW     CM - FL + DGB CM - FL + DGB + PCW     CM - FL + DGB CM - FL + DGB + PCW Centralized + GB (Upperbound)
0 $\simeq$ 36.52 $\simeq$ 36.52     $\simeq$ 36.52 $\simeq$ 36.52     $\simeq$ 36.52 $\simeq$ 36.52 $\simeq$ 36.52
100 $\simeq$ 56.56 $\simeq$ 57.07     $\simeq$ 46.26 $\simeq$ 49.93     $\simeq$ 46.83 $\simeq$ 55.32 $\simeq$ 58.59
200 $\simeq$ 65.65 $\simeq$ 66.67     $\simeq$ 48.25 $\simeq$ 58.63     $\simeq$ 53.82 $\simeq$ 60.82 $\simeq$ 69.70
300 $\simeq$ 67.67 $\simeq$ 68.69     $\simeq$ 49.63 $\simeq$ 64.32     $\simeq$ 55.72 $\simeq$ 64.17 $\simeq$ 74.24
400 $\simeq$ 69.19 $\simeq$ 70.20     $\simeq$ 52.36 $\simeq$ 66.24     $\simeq$ 51.93 $\simeq$ 66.82 $\simeq$ 77.27
500 $\simeq$ 70.71 $\simeq$ 71.21     $\simeq$ 53.64 $\simeq$ 67.35     $\simeq$ 56.72 $\simeq$ 68.39 $\simeq$ 78.28
600 $\simeq$ 71.72 $\simeq$ 72.22     $\simeq$ 53.83 $\simeq$ 67.84     $\simeq$ 57.26 $\simeq$ 69.92 $\simeq$ 78.79
TABLE XVII: The best accuracy (%) acquired by all methods alongside the gap with our method across different data sizes used per institution for binary and multi-class classification under type-based heterogeneity.
Binary Classification Multi-Class Classification
Dataset Size Method CM-FL HGB-FL CM-FL + DGB Our Method Dataset Size Method CM-FL HGB-FL CM-FL + DGB Our Method
30 $\simeq$ 60.48 $\simeq$ 64.52 $\simeq$ 66.12 $\simeq$ 70.96 30 $\simeq$ 47.97 $\simeq$ 51.51 $\simeq$ 47.97 $\simeq$ 63.13
40 $\simeq$ 63.71 $\simeq$ 66.93 $\simeq$ 68.54 $\simeq$ 72.58 40 $\simeq$ 48.48 $\simeq$ 54.04 $\simeq$ 50.51 $\simeq$ 65.66
50 $\simeq$ 64.04 $\simeq$ 68.64 $\simeq$ 69.21 $\simeq$ 74.16 50 $\simeq$ 49.49 $\simeq$ 57.52 $\simeq$ 53.83 $\simeq$ 67.84
60 $\simeq$ 64.52 $\simeq$ 75.06 $\simeq$ 76.61 $\simeq$ 78.22 60 $\simeq$ 50.51 $\simeq$ 62.63 $\simeq$ 59.60 $\simeq$ 71.71
TABLE XVIII: The best accuracy (%) acquired by all methods alongside the gap with our method across different data sizes used per institution for binary and multi-class classification under class-based heterogeneity.
Binary Classification Multi-Class Classification
Dataset Size Method CM-FL HGB-FL CM-FL + DGB Our Method Dataset Size Method CM-FL HGB-FL CM-FL + DGB Our Method
30 $\simeq$ 61.29 $\simeq$ 66.93 $\simeq$ 67.74 $\simeq$ 72.58 30 $\simeq$ 47.98 $\simeq$ 52.02 $\simeq$ 49.49 $\simeq$ 64.65
40 $\simeq$ 62.03 $\simeq$ 68.65 $\simeq$ 70.16 $\simeq$ 75.02 40 $\simeq$ 48.48 $\simeq$ 55.56 $\simeq$ 51.52 $\simeq$ 67.17
50 $\simeq$ 65.24 $\simeq$ 69.15 $\simeq$ 71.13 $\simeq$ 76.31 50 $\simeq$ 48.53 $\simeq$ 63.56 $\simeq$ 57.26 $\simeq$ 69.92
60 $\simeq$ 66.12 $\simeq$ 77.41 $\simeq$ 75.80 $\simeq$ 79.03 60 $\simeq$ 52.02 $\simeq$ 64.14 $\simeq$ 61.62 $\simeq$ 73.23
We compare the values computed during the experiments on each of the heterogeneity types, i.e., the type-based and the class-based heterogeneities, for both binary-class and multi-class cases. In particular, the bar plots are obtained via observing the statistical distribution of the values of $\overline{\rho}_{C,n}^{(t)}$ over the training time $t$ for each institution $n$ (x-axis label of the figures) containing modality combination $C$ (represented in the legend of the figures). Focusing on type-based heterogeneity, given the distributions of cohort data in Table IV, the plots in Fig. 5(a) and 5(b) show the values of $\overline{\rho}_{C,n}^{(t)}$ for each institution $n\in\mathcal{N}$, where the numbers above each bar in the plots show the institutions corresponding Category in Table IV. As can be deducted from Table IV, datasets of institutions belonging to Category 1, closely resembles the global distribution of data in Table II. [^12] This is obtained by comparing the percentage of data points belong to each cohort in each institution in Table IV to that of the global dataset in Table II. Also, it can be seen that the closeness of the local data distribution to that of the global one drops across Category 2 and 3. Thus, intuitively, we expect that, for each modality combination, institutions belonging to Category 1 contribute the most to the calculation of DOGR parameters, followed by institutions in the Categories 2 and 3. This is in fact what the plots in Fig. 5(a) and 5(b) show: for each modality combination, shown via a unified color, institutions from Category 1 have the highest values of $\rho$ followed by those belonging to Category 2 and 3. Further, since the order of the values of $\rho$ mimics the same order in data heterogeneity of the institutions, the results also validate that our proposed metric in PCW can differentiate between the levels of similarity of local datasets of institutions. The same deduction can be made focusing on class-based heterogeneity via inspecting the results depicted in Fig. 5(c) and 5(d), noting that the categorization of institutions in Table V is similar to Table IV: Category 1 closely resembles the global distribution of data in Table II and the closeness of the local data distribution to that of the global one drops across Category 2 and 3.
V-E5 Comparing the performance of the trained models across various cancer types
Given the co-existence of multiple cancer types in the dataset considered, we also study how the trained global model performs when applied on different cancer types in isolation. A higher difference in the test accuracy of the global model on the data of multiple cohorts indicates may further imply a more unbalanced training. In order to measure the performance, we have taken the global model with the best validation accuracy during the training for each method, and tested it against the validation data of each cohort. The results for each case are included in Table XIII and XIV. It is shown that in the majority of cases, the performance of baseline methods exhibit large variations across the cohorts. In particular, in most of the cases, the performance of the baseline methods are skewed towards the BRCA cohort. This can indicate the effect of data imbalance between BRCA and the other two existing cohorts (see Table II) shifts the learning parameters toward the BRCA data characteristics and therefore biasing the global model and parameters to become more suitable for breast cancer classification. Also, it can be seen that, our method exhibits the best performance across all the individual cohorts in both Table XIII and XIV. Further, although our method was not directly developed to improve the ‘fairness’ across various cancer types, it still reduces the gap between the performance across various cohorts, indicating a more balanced learning. This further opens a new avenue of research on the development of novel techniques for multi-modal FL in the cancer staging context, which are focused to improve the learning fairness across various cancer types.
V-E6 Comparing the performance of the trained models under iid and non-iid data distribution across the institutions
In order to observe the effect of PCW in addressing data heterogeneity, we conduct an experiment where we compare the performance of the global model under homogeneous/iid data and heterogeneous/non-iid distributions. In the iid setting, we have sampled the data uniformly at random to create each institution’s local dataset. In the non-iid case, we have implemented both type-based and class-based heterogeneities. The results of our experiments for both binary and multi-class classification cases are shown in Table XV and XVI. As our results show, the implementation of DGB in an iid scenario (second column) reaches a higher accuracy from the implementation of DGB in a non-iid scenario (fourth and sixth columns) espeically as the training progresses; thus proving that non-iid-ness induces a notable performance drop in distributed/FL setups. The addition of PCW in a non-iid scenario (fifth and seventh columns) observably elevates the performance reached when compared with the same-type heterogeneity when PCW is not applied (fourth and sixth columns). This also proves that PCW addresses the data heterogeneity issue in the distributed/FL setups, given that the only difference between the two scenarios is the utilization of PCW. The last column in each table also gives the result for the centralized scenario where gradient blending is applied, showing an upper bound on the performance where the heterogeneity of data is not present. Comparing the centralized case (last column) and the non-iid cases where PCW is not used (fourth and sixth columns), we can see how much the accuracy is degraded due to the heterogeneity of data (at most $16.27$% in the case of binary classification in Table XV and $24.96$% in the case of multi-class classification in Table XVI). Comparing the results where PCW has been used in non-iid scenarios (fifth and seventh columns of these two tables) with the case that we have not used PCW (fourth and sixth columns), we find that PCW through addressing the presence of non-iid data has decreased the degradation in performance when compared to centralized training as follows:
  • Binary Classification, Type-based Heterogeneity: The gap is decreased by $4.95\%=74.16\%-69.21\%$. In particular, the performance gap to centralized training after applying the PCW is $11.32\%=85.48\%-74.16\%$ while this gap is $16.27\%=85.48\%-69.21\%$ when PCW is not applied.
  • Binary Classification, Class-based Heterogeneity: The gap is decreased by $5.18\%=76.31\%-71.13\%$. In particular, the performance gap to centralized training after applying the PCW is $9.17\%=85.48\%-76.31\%$ while this gap is $14.35\%=85.48\%-71.13\%$ when PCW is not applied.
  • Multi-class Classification, Type-based Heterogeneity: The gap is decreased by $14.01\%=67.84\%-53.83\%$. In particular, the performance gap to centralized training after applying the PCW is $10.95\%=78.79\%-67.84\%$ while this gap is $24.96\%=78.79\%-53.83\%$ when PCW is not applied.
  • Multi-class Classification, Class-based Heterogeneity: The gap is decreased by $12.66\%=69.92\%-57.26\%$. In particular, the performance gap to centralized training after applying the PCW is $8.87\%=78.79\%-69.92\%$ while this gap is $21.53\%=78.79\%-57.26\%$ when PCW is not applied.
V-E7 Comparing the performance of different methods across various data availability scenarios
In order to gain a better understanding of the level of performance degradation that the scarcity of data imposes on each method, we conduct experiments under different numbers of datapoints available per institution. [^13] Note that due to the size of the overall dataset, the maximum number of available datapoints per institution is $60$, which is reflected in Tables XVII and XVIII. The results are presented in Tables XVII and XVIII where the experiments have been conducted for the type-based and class-based heterogeneity (the results are obtained after $600$ global aggregation steps). It can be seen that, our method, which explicitly aims to address the heterogeneity of data (both at the modality combinations level and data belonging to each class label/cancer type available across the institutions), has been able to result in a superior performance regardless of the number of datapoints available at the institutions.
It is worth mentioning that this case study further opens a research direction on the study and adaptation of methods that explicitly aim to address the scarcity of data (e.g., few shot learning [39]) to multi-modal FL under unbalanced presence of modalities across the participating institutions, an area which is yet to be explored.